OpenStudy (anonymous):

Can someone check my answer please? Petra deposited $278.16 in a savings account that earns 2.1% simple interest. What is Petra’s account balance after nine years? I'm stuck between $52.57 and $330.73

OpenStudy (jdoe0001):

Interest = principal(deposit) * rate(in decimals) * time( in years) i = prt deposit = $278.16 rate = 2.1% in decimals will be \(\bf \cfrac{2.1}{100}=0.021\) time = 9 years Petra's account will be, whatever she deposited, PLUS the interest \(\bf \color{red}{i} = prt\\ \quad \\ \color{red}{i} = 278.16\times 0.021\times 9\\ \quad \\ \textit{her balance after 9 years is}\qquad 278.16 + \color{red}{i} \)

OpenStudy (anonymous):

I just reread the question and$52.57 wouldn't be correct at all. But I got $330.73 by using I = Prt +P. Is that what you were saying? Because I did use I = Prt but I ended up with $52.57 which is too low

OpenStudy (anonymous):

So now I'm just trying to make sure $330.73 is the correct answer

OpenStudy (jdoe0001):

52.57 is correct for the INTEREST earned only

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Got it, and 330.73 would be her ACCOUNT balance after 9 years. Right? (:

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Sweet! Thank you very much for checking! :D

OpenStudy (jdoe0001):

yw